Assume the following ratios are constant.

Total asset turnover = 2.24
Profit margin = 5.2 %
Equity multiplier = 1.71
Payout ratio = 49 %

What is the sustainable growth rate? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

Sustainable growth rate             %
Answer #1

ROE=Profit margin*Total Asset Turnover*Equity multiplier

=(0.052*2.24*1.71)=0.1991808

Retention ratio=1-payout ratio

=(1-0.49)=0.51

Sustainable Growth Rate=(ROE*Retention ratio)/[1-(ROE*Retention ratio)]

=(0.1991808*0.51)/[1-(0.1991808*0.51)]

which is equal to

=11.31%(Approx).